OpenStudy (anonymous):

Which method can help me to do the integral of (x^2+1)/(1-x^2)?

OpenStudy (aravindg):

quotient rule may be

OpenStudy (nenadmatematika):

write x^2+1 as x^2-1+2, make two basic integrals and solve

OpenStudy (anonymous):

Because I just learned in school to do a partial fraction decomposition but this somehow does not work out very well. I´ll use your tip nenadmatematika.

OpenStudy (dumbcow):

ha i was just going to suggest partial fractions 1-x^2 = (1-x)(1+x)

OpenStudy (dumbcow):

oh right split it into 2 fractions first

OpenStudy (nenadmatematika):

\[=\int\limits_{?}^{?}(x^2-1)/(1-x^2)dx+\int\limits_{?}^{?}2/(1-x^2)=-\int\limits_{?}^{?}dx+2\int\limits_{}^{?}dx/(1-x^2)\]\[=-x+\ln \left| (1+x)/(1-x) \right|+c\]

OpenStudy (anonymous):

This kind of idea comes from experience I guess?

OpenStudy (nenadmatematika):

yes, it takes some practice to get the routine, but consider it very easy integral...:D

OpenStudy (anonymous):

ok, thx for your help :)

OpenStudy (nenadmatematika):

if you've got some more problems with integrals hit me now :D

OpenStudy (turingtest):

Do you want to see it with partial fractions? it's not very hard

OpenStudy (anonymous):

Yes, might be a nice to see a different approach.

OpenStudy (turingtest):

well actually, I guess nenad did use them it do the second integral (or at least I don't see how he did without them) another way to get to that point it to rearrange the fraction so you can perform long division\[\frac{1+x^2}{1-x^2}=\frac{-x^2-1}{x^2-1}=-1-\frac2{x^2-1}\]now doing partial fractions on the second term\[\frac{-2}{x^2-1}=\frac A{x-1}+\frac B{x+1}\]\[A(x+1)+B(x-1)=-2\]\[A=-1,B=1\]so we wind up with\[\int-1+\frac1{x-1}-\frac1{x+1}dx\]again

OpenStudy (nenadmatematika):

well, that's nice but you would you make that second part complicated when you know that it's a basic integral: \[\int\limits_{?}^{?}dx/(1-x^2)=(1/2)\ln \left| (1+x)/(1-x)\right|\]

OpenStudy (turingtest):

I didn't know that form, but if they are supposed to use partial fractions in the problem they must do it here.

OpenStudy (nenadmatematika):

OK, I didn't know that they must do it like that :D

OpenStudy (anonymous):

Can I do this partial fractions stuff also without changing x^2+1 to x^2-1+2?

OpenStudy (turingtest):

the order of the numerator has to be smaller than that in the denominator to do partial fractions, so you need to make that happen somehow long division is a very standard procedure in these cases, but you can also do what nenad did frequently.

OpenStudy (turingtest):

the degree of the numerator*

OpenStudy (anonymous):

Ah, Thank you. We only talked in school about the cases where the degree is higher or lower but never about what it means when they have the same degree. :( My teacher sucks... Thank you TuringTest for your help and explanation.